Find the number which when decreased by three-fourths of itself equals 28.

Answer by checkley79(3341)

X-3X/4=28
X/4=28
X=4*28
X=112 ANS.
PROOF:
112-3*112/4=28
112-336/4=28
112-84=28
28=28